. A certain hydrocarbon on complete combustion at s.t.p produced 89.6dm3 of CO2
and 54g of water. The hydrocarbon should be
1
Expert's answer
2020-02-25T04:53:42-0500

moles of CO2 ="\frac{V}{V_m}=\frac{89.6}{22.4} = 4 mole"

"n(C) = n(CO_2)=4 mole"

moleas of H2O ="\frac{m}{M} =\frac{54}{18} = 3 mole"

"n(H) = 2\times n(H_2O)=2\times 3 = 6 mole"

mole ratio "n(C);n(H) = 4:6= 2:3"

the empirical formula of hydrocarbon is C2H3. the molecular formulas could be:

C4H6